# Semantic Web – A Forgotten Wave of Artificial Intelligence?

### v1

@@ Abstract
Artificial Intelligence, Software Agents, Semantic Web, World Wide Web, Large Language Models
@@ 1 Introduction
The notion of Artificial Intelligence (AI) as a distinct discipline within Computer Science was introduced in 1956 at a workshop at Dartmouth College by early pioneers such as John McCarthy, Marvin Minsky, and others. Since then, the history of AI has experienced recurring cycles of enthusiasm and subsequent disillusionment, commonly referred to as “AI summers” and “AI winters” [36]. These cycles are associated with two main approaches to creating AI: symbolic (“white box”) approaches, based on explicit knowledge representation, logic, reasoning, and search, and sub-symbolic (“black box”) approaches, based on neural networks, machine learning, and statistical models.
@@ 2.1 The Hype Cycle
The second equation, the S-curve of technological maturity, represents the gradual advancement of a technology over time. Initially, progress is slow due to limited understanding and minimal performance gains from early investments. However, as knowledge accumulates and improvements accelerate, the technology experiences rapid development until it eventually stabilizes at a plateau defined by its inherent limitations. Together, these two curves form the hype cycle (solid gray line), which models the typical trajectory of emerging technologies, highlighting both the initial overhype and the eventual stabilization at a realistic level of maturity.
The stages outlined in the hype cycle depicted in Figure 2 are closely related to the way many studies have seen the historical trajectory of AI:
@@
These stages, which reflect periods of intense optimism and disillusionment, are commonly referred to as “AI summers” and “AI winters”. During AI summers, expectations for AI technologies reach the “Peak of Inflated Expectations”, driven by breakthroughs and media enthusiasm, while subsequent AI winters align with the “Trough of Disillusionment”, as the technology fails to meet overinflated promises. Over time, as understanding deepens and practical applications emerge, AI progresses through the “Slope of Enlightenment” to a stable “Plateau of Productivity”, marking phases of realistic growth and adoption.
2.2 Review and Synthesis of AI History Illustrations
In this section, we present illustrations of AI history. The insights and illustrations discussed here are derived from a combination of sources, including scientific papers, blog posts, and public presentations available on the internet, providing a comprehensive perspective on the cyclical nature of AI development.
@@ 2.3 Summary of AI Illustrations
All AI history illustrations known to the authors of this study depict the period from 2000 to 2010 as an AI winter rather than a time of substantial research advancements. These accounts consistently emphasize stagnation and minimal innovation, portraying this period as one characterized by diminished funding, waning public interest, and a general lack of transformative breakthroughs in AI. While some research activities continued during this time, the prevailing perception is that AI did not experience any major surge or revolution.
@@ 3 The Semantic Web and Its Vision of AI Software Agents
3.1 Software Agents Before Semantic Web
@@
The development of various agent architectures, such as reactive, deliberative, and hybrid models, further illustrated the field’s commitment to creating sophisticated autonomous entities [62]. Researchers also focused on defining the boundaries of agency, distinguishing agents from conventional programs through taxonomies and formal definitions [15]. The intellectual roots of software agents were traced back to Distributed Artificial Intelligence (DAI) and Carl Hewitt’s Actor model from 1977, highlighting the long-standing interest in distributed and interacting intelligent systems [45, 19, 64]. This period laid a solid foundation for the subsequent exploration of synergies between software agents and emerging technologies like the Semantic Web.
@@ 3.2 The Vision of the Semantic Web and Intelligent Agents
Tim Berners-Lee, James Hendler, and Ora Lassila, in their seminal 2001 paper ”The Semantic Web” [5], envisioned an evolution of the World Wide Web into a more intelligent and interconnected system. This Semantic Web would be an infrastructure where information has explicitly defined meaning, facilitating human-computer cooperation and enabling machines to understand and process web data through a universal framework of metadata and ontologies. This involved enriching web resources with semantic markup using languages like RDF and OWL, making content machine-processable [37]. The goal was a transition from unstructured documents to interconnected data, enhancing knowledge sharing, reuse, and integration [53], aligning directly with AI’s goals of machine understanding and reasoning [5]. As of the time of writing, this highly influential paper has garnered over 30,000 citations.
@@ 3.3 Key Enabling Technologies and the Semantic Web Stack
Agent Communication Languages (ACLs), such as the Foundation for Intelligent Physical Agents (FIPA) ACL [12], were recognized as essential for facilitating interaction and information exchange between autonomous agents. Furthermore, initiatives like OWL-S (later known as DAML-S) [35] aimed to provide semantic descriptions of web services, allowing agents to automatically discover, invoke, and compose these services based on their semantic properties. The development of these technologies demonstrates a focused effort to build the necessary infrastructure for intelligent agents to function effectively within the Semantic Web, with a strong emphasis on knowledge representation (ontologies, RDF, OWL) and inter-agent communication (ACLs, Semantic Web Services).
Figure 4 illustrates the ”Semantic Web Stack,” a layered architecture proposed by Tim Berners-Lee for organizing the technologies of machine-readable web data [4]. URIs and Unicode provide the foundation for identification and character encoding. XML offers a syntax for structured documents, while RDF and RDF Schema (RDFS) define a standardized way to describe resources and their relationships [7]. OWL builds on this to enable the creation of complex ontologies. Rule languages like RIF (Rule Interchange Format) and SWRL (Semantic Web Rule Language), along with the SPARQL query language, were intended to support inference and querying over this semantically rich data [22]. Higher layers of the stack were envisioned to handle logical reasoning, proof verification, and trust establishment through cryptographic methods, ultimately enabling applications with enhanced interoperability.
@@ 3.4 Software Agents in the LLM Age 2017 onwards
The advent of large language models (LLMs) has dramatically reshaped the landscape of software agent development, opening up new possibilities and challenges. Pre-LLM, agents often relied on meticulously crafted rules, ontologies, and knowledge bases, limiting their adaptability and generalizability [64, 51]. While these approaches provided a degree of control and explainability, they struggled to handle the complexity and ambiguity of real-world scenarios. LLMs, trained on massive datasets, offer a fundamentally different approach. They possess an unprecedented ability to understand and generate natural language, reason, and even exhibit some degree of common sense, albeit imperfectly [8, 46, 60]. This inherent capability allows LLM-powered agents to interact with users and environments in a much more natural and flexible way, reducing the need for extensive manual programming and domain-specific knowledge engineering.
The integration of LLMs with agent frameworks has led to the emergence of ”LLM-powered agents” or ”LLM agents,” capable of performing a wider range of tasks than their predecessors. These agents can leverage LLMs for various functions, including natural language understanding (NLU), natural language generation (NLG), planning, decision-making, and even tool use [67, 59, 68]. For instance, an LLM agent can be prompted to break down a complex task into smaller, manageable sub-tasks, search the web for relevant information, interact with APIs to execute actions, and summarize the results for the user [69, 54]. Frameworks like LangChain [9] and AutoGen [66] provide tools and abstractions to facilitate the development and deployment of these agents, enabling developers to chain together LLM calls, external tools, and custom logic. This marks a significant shift from symbolic AI to a more data-driven, emergent approach to agent development.
@@
The ongoing research in LLM-powered agents focuses on addressing these limitations and exploring new capabilities. This includes developing methods for improving the factual accuracy and reliability of LLMs, designing more robust and explainable agent architectures, and creating effective evaluation metrics to assess agent performance across diverse tasks [43, 48]. The convergence of LLMs and software agents represents a significant step towards building more general-purpose, adaptable, and intelligent systems, with the potential to transform how we interact with computers and automate complex tasks. However, responsible development and deployment, with a keen awareness of the inherent challenges, are crucial to realizing the full potential of this technology.
For example, Microsoft CEO Satya Nadella recently stated that AI-driven Software Agents are fully revolutionizing business-to-business (B2B) applications. According to Nadella, AI-driven agents will take over business logic and enable new AI-powered applications, placing software agents at the heart of the current AI hype. In December 2024, during an episode of the BG2 Pod series, Nadella stated in the interview video [42], e.g., that:
  • “I think that the notion that business applications (B2B Apps) exist, that is probably where they all collapse, in the (Software) Agent-era.”
  • “Because if you think about it, they (B2B Apps) are essentially CRUD-databases with a bunch of business logic. The business logic will go to these Agents. These Agents are going to be multi-repo-CRUD so they are not going to discriminate between what the backend is. They are going to be updating multiple databases. All the (business) logic will be in the AI-tier (Agent-layer) so to speak.”
  • “Once the AI-tier becomes the place where all the logic is, then people will start replacing the backends.”
  • “As we speak, I think we are seeing quite high rates of wins in our Dynamics backends and the Agent use. We will go very aggressive trying to collapse it all. Whether it is customer service, be it finance and operations…”
  • “People want more AI-native business apps. It means the logic tier can be orchestrated by AI and AI Agents in a very seamless way.”

### v2

# Semantic Web And Software Agents – A Forgotten Wave of Artificial Intelligence?

@@ Abstract
Artificial Intelligence, Software Agents, Semantic Web, World Wide Web, Hypertext
@@ 1 Introduction
The notion of Artificial Intelligence (AI) as a distinct discipline within Computer Science was introduced in 1956 at a workshop at Dartmouth College by early pioneers such as John McCarthy, Marvin Minsky, and others. Since then, the history of AI has experienced recurring cycles of enthusiasm and subsequent disillusionment, commonly referred to as “AI summers” and “AI winters” [47]. An AI winter refers to a period of reduced funding, interest, and progress in AI research, typically following unmet expectations or technological limitations. In contrast, an AI summer denotes a phase of heightened enthusiasm, investment, and innovation, often driven by breakthroughs in algorithms, computing power, or applications. These cycles are associated with two main approaches to creating AI: symbolic (“white box”) approaches, based on explicit knowledge representation, logic, reasoning, and search, and sub-symbolic (“black box”) approaches, based on neural networks, machine learning, and statistical models.
@@ 2.1 The Hype Cycle
The second equation, the S-curve of technological maturity, represents the gradual advancement of a technology over time. Initially, progress is slow due to limited understanding and minimal performance gains from early investments. However, as knowledge accumulates and improvements accelerate, the technology experiences rapid development until it eventually stabilizes at a plateau defined by its inherent limitations. Together, these two curves form the hype cycle (solid gray line), which models the typical trajectory of emerging technologies, highlighting both the initial overhype and the eventual stabilization at a realistic level of maturity. The stages outlined in the hype cycle depicted in Figure 2 are:
@@
Although the hype-cycle is typically illustrated as a single expectation curve, it actually overlays the S-curve of technology diffusion; how these two curves interact determines both the tempo and the amplitude of successive boom-and-bust phases, yet this relationship remains only sketchily articulated in the literature. Consequently, existing research offers limited guidance on when firms and investors should initiate, scale, or exit their involvement in order to capture upside while containing downside risk. Closing this gap will require analytical frameworks that fuse expectation dynamics with adoption kinetics and strategic positioning along the hype-cycle.
2.2 Review and Synthesis of AI Winters and AI Summers
In this section, we synthesise evidence from peer-reviewed papers, technical blogs, and public presentations to trace the recurring boom-and-bust pattern (AI Summers and AI Winters) that has characterised artificial intelligence research and commercialisation. It is important to note that AI is not a single invention, but a shifting portfolio of techniques that is constantly redefined over time — from symbolic reasoning and expert systems to machine learning, deep learning, and today’s large-scale generative models. Each element in this portfolio follows its own trajectory, so the broad history of AI is best understood as the superposition of many overlapping hype cycles rather than one monolithic wave.
These technology-specific curves explain why the field repeatedly climbs to a Peak of Inflated Expectations, descends into a Trough of Disillusionment, and, where genuine utility is present, climbs the Slope of Enlightenment to a durable Plateau of Productivity. Recognising AI as a constellation of evolving technologies helps clarify why optimism and disappointment can coexist—and why practical value often emerges only after multiple passes through the cycle.
Take neural networks and software agents as illustrative examples. For a long time, neural networks were not considered mainstream AI [27]; however, especially during the past 15 years, that has changed. Similarly, introduced in the early 1990s, agent technology ascended to widespread media enthusiasm, encountered technical and market barriers, resurfaced with the rise of web services, and is now experiencing renewed interest through autonomous AI assistants. In other words, the same underlying concept has already traversed the Gartner hype cycle several times.
@@ 2.3 Summary of AI Illustrations
AI history illustrations known to the authors of this study depict the period from 1995 to 2010 as an AI winter rather than a time of substantial research advancements. These accounts consistently emphasize stagnation and minimal innovation, portraying this period as one characterized by diminished funding, waning public interest, and a general lack of transformative breakthroughs in AI. While some research activities continued during this time, the prevailing perception is that AI did not experience any major surge or revolution.
@@ 3 The Semantic Web and Its Vision of AI Software Agents
3.1 Hype Cycles: Software Agent Technology
Software agent technology has gone through a hype cycle a couple of times during its history. The following quote is from 1996:
”Agents are here to stay, not least because of their diversity, their wide range of applicability and the broad spectrum of companies investing in them. As we move further and further into the information age, any information-based organization which does not invest in agent technology may be committing commercial hara-kiri.”  [58]
Authors of the above quote wrote another article a few years later  [59] stating that the assumptions were too high. If anything, companies that invested in agent technology in the 1990s were perhaps the ones to commit a form of hara-kiri.
Recently, in December 2024, Microsoft CEO Satya Nadella stated that AI-driven software agents are fully revolutionizing business-to-business (B2B) applications. According to Nadella, AI-driven agents will take over business logic and enable new AI-powered applications, placing software agents at the heart of the current AI hype. The full quote from an episode of the BG2 Pod series, where Nadella stated in an interview video [53], e.g., is:
“I think that the notion that business applications (B2B Apps) exist, that is probably where they all collapse, in the (Software) Agent-era.” – “Because if you think about it, they (B2B Apps) are essentially CRUD-databases with a bunch of business logic. The business logic will go to these Agents. These Agents are going to be multi-repo-CRUD so they are not going to discriminate between what the backend is. They are going to be updating multiple databases. All the (business) logic will be in the AI-tier (Agent-layer) so to speak.” “Once the AI-tier becomes the place where all the logic is, then people will start replacing the backends. – As we speak, I think we are seeing quite high rates of wins in our Dynamics backends and the Agent use. We will go very aggressive trying to collapse it all. Whether it is customer service, be it finance and operations…” – “People want more AI-native business apps. It means the logic tier can be orchestrated by AI and AI Agents in a very seamless way.”
Obviously, noting that agent technology has had inflated expectations in the past does not prove anything about the future. This time, it might be different, as the overall technology infrastructure has evolved significantly since the 1990s. We also have to remember that, for example, neural networks have experienced several hype cycles before the major breakthrough in the 2010s.
3.2 Software Agents Before the Semantic Web
Long before the Web era, the conceptual foundations for software agents was established by research in Distributed Artificial Intelligence (DAI) and the Actor model [26]. DAI cast intelligence as an emergent phenomenon arising from multiple, loosely-coupled problem solvers that cooperate across a network, thereby elevating autonomy and negotiation to primary design concerns [8, 22]. In parallel, the Actor model supplied a concrete computational metaphor of independent, concurrently executing entities that communicate solely through asynchronous message passing. Together these two traditions contributed the twin conceptual pillars of distribution and concurrency, which later researchers distilled into the now-standard agent properties of social ability and proactive–reactive behaviour [81].
@@ 3.2 Software Agents Before the Semantic Web
The development of various agent architectures, such as reactive, deliberative, and hybrid models, further illustrated the field’s commitment to creating sophisticated autonomous entities [79]. Researchers also focused on defining the boundaries of agency, distinguishing agents from conventional programs through taxonomies and formal definitions [21]. The intellectual roots of software agents were traced back to Distributed Artificial Intelligence (DAI) and Carl Hewitt’s Actor model from 1977, highlighting the long-standing interest in distributed and interacting intelligent systems [58, 26, 81].
From Definitions to Standards.
The explicit taxonomies and formal definitions produced during this period played a pivotal role in standardising the concept of agency. By establishing a common vocabulary, they enabled the Foundation for Intelligent Physical Agents (FIPA) to publish its initial specifications—most notably the FIPA Agent Communication Language (FIPA ACL) and a reference architecture that has since underpinned almost every industrial-grade multi-agent platform [18]. These efforts, together with subsequent W3C initiatives (e.g. the Agent Markup Language submission), ensured that later Semantic-Web agents could interoperate across organisational and domain boundaries on the basis of well-defined semantics rather than ad-hoc conventions.
3.2.1 Impact of the World Wide Web’s Expansion on Software Agents
The exponential expansion of the World Wide Web in the mid-1990s profoundly altered the landscape in which software agents were conceived and applied. Originally developed in the context of distributed AI and local computation environments, software agents found new significance as the Web became a vast, decentralized, and heterogeneous information space. The growing complexity of navigating, searching, and processing this information catalyzed the demand for autonomous systems capable of filtering, retrieving, and reasoning over web-based content on behalf of users.
Seminal work by Pattie Maes and others [45] recognized this shift early, introducing agents as personalized assistants that could reduce information overload in the emerging Web era. This period also saw the rise of ”information agents”—autonomous programs designed specifically to operate over Web protocols, harvest content, and interact with web services. The scalability and openness of the Web provided a compelling environment for agents to demonstrate their core properties: autonomy, proactivity, and social interaction, especially as HTML and HTTP standards facilitated agent-based crawling and retrieval.
Moreover, the ubiquity of browsers and online services created a new user expectation: software should act intelligently in navigating and managing web interactions. This sociotechnical transformation fueled both academic research and industrial exploration into agent-based architectures, eventually motivating the integration of agents into the foundational vision of the Semantic Web [6]. Thus, the Web’s growth did not merely provide a backdrop—it actively shaped the functionality, relevance, and perceived utility of software agents during this foundational decade.
@@ 3.3 The Vision of the Semantic Web and Intelligent Agents
Tim Berners-Lee, James Hendler, and Ora Lassila, in their seminal 2001 paper “The Semantic Web” [6], envisioned an evolution of the World Wide Web into a more intelligent and interconnected system. This Semantic Web would be an infrastructure where information has explicitly defined meaning, facilitating human-computer cooperation and enabling machines to understand and process web data through a universal framework of metadata and ontologies. This involved enriching web resources with semantic markup using languages like RDF and OWL, making content machine-processable [48]. The goal was a transition from unstructured documents to interconnected data, enhancing knowledge sharing, reuse, and integration [66], aligning directly with AI’s goals of machine understanding and reasoning [6]. As of the time of writing, this highly influential paper has garnered over 30 000 citations.
@@
Realising this vision depends on the ability of autonomous agents to converse. The FIPA ACL specification meets this requirement by supplying a performative-based message envelope whose semantics are defined in a modal logic of communicative acts [19]. When the content field of such messages is expressed in RDF or OWL, the resulting utterances are simultaneously interpretable at the data and the speech-act levels, thereby enabling richly contextualised dialogues among heterogeneous Web agents [75]. Accordingly, ACLs form the critical bridge between the Semantic Web’s data layer and the multi-agent community’s interaction layer.
@@ 3.4 Key Enabling Technologies and the Semantic Web Stack
Agent Communication Languages (ACLs), such as the Foundation for Intelligent Physical Agents (FIPA) ACL [16], were recognized as essential for facilitating interaction and information exchange between autonomous agents. Furthermore, the DAML-S ontology—renamed OWL-S after the release of OWL—[46] was created to add a machine-interpretable semantic layer to conventional Web-service descriptions. It exposes three mutually supportive components: a Service Profile that advertises what the service does, a Process Model that specifies how the service can be executed, and a Grounding that binds abstract inputs and outputs to concrete message exchanges. Together these components allow an intelligent agent to match its goals to advertised profiles and thus discover relevant services, to use the process model and grounding information to construct and send the correct messages for direct invocation, and to chain multiple process models into coherent workflows for automatic composition of complex tasks. The development of DAML-S/OWL-S therefore represents a focused effort to supply the semantic infrastructure—rooted in ontologies, RDF, and OWL—along with the inter-agent communication mechanisms required for truly autonomous agents on the Semantic Web.
Figure 4 illustrates the ”Semantic Web Stack,” a layered architecture proposed by Tim Berners-Lee for organizing the technologies of machine-readable web data [5]. URIs and Unicode provide the foundation for identification and character encoding. XML offers a syntax for structured documents, while RDF and RDF Schema (RDFS) define a standardized way to describe resources and their relationships [9]. OWL builds on this to enable the creation of complex ontologies. Rule languages like RIF (Rule Interchange Format) and SWRL (Semantic Web Rule Language), along with the SPARQL query language, were intended to support inference and querying over this semantically rich data [30].
While OWL enables inference through description logic constructs such as subclassing, property hierarchies, and transitive closures, its expressivity is intentionally constrained to ensure decidability and tractable reasoning. To address use cases that require more expressive logic—such as conditional rules and inter-property dependencies—rule-based extensions were introduced. The Semantic Web Rule Language (SWRL) [30] allows the combination of OWL ontologies with Horn-like rules of the form if antecedent then consequent, significantly enhancing inferencing capabilities. For example, SWRL can express rules like ”if a person has a parent who is female, then the parent is a moth
Higher layers of the stack were envisioned to handle logical reasoning, proof verification, and trust establishment through cryptographic methods, ultimately enabling applications with enhanced interoperability. Over the past two decades, the Semantic Web community has converged on a set of concrete cryptographic primitives that operationalise the previously aspirational proof and trust layers. A prerequisite for any digital signature is a deterministic byte-level representation of the graph being protected; this is now provided by the RDF Dataset Canonicalisation and Hash 1.0 (RDFC-1.0) algorithm, which yields a stable hash despite blank-node renaming and graph isomorphisms [43]. Once canonicalised, graphs can be enveloped by the Data Integrity 1.0 suite (formerly “Linked-Data Signatures”), embedding cryptographic proofs directly in RDF or JSON-LD and supporting selective-disclosure cryptosuites such as BBS+ [71].
Building on these primitives, the Verifiable Credentials Data Model 2.0 binds claims to subjects in tamper-evident envelopes that agents can verify offline, while Decentralized Identifiers (DID) Core replaces hierarchical certificate authorities with URL-resolvable public-key documents [70, 69]. Together these technologies allow autonomous agents—whether classic rule engines or modern LLM pipelines—to (i) verify the provenance of data before reasoning, (ii) attach signed proofs to their own outputs, and (iii) assemble verifiable chains of custody across heterogeneous knowledge graphs.
@@ 3.4.1 Semantic Web Rollout Vision
3.5 Semantic Web Influence on Modern AI/ML Data Integration
Although the Semantic Web’s most ambitious goals remain aspirational, its foundational interoperability standards—such as RDF, OWL, SPARQL, and SHACL—have significantly influenced how structured data is integrated in modern AI and machine learning systems. These standards provided early blueprints for achieving semantic consistency across heterogeneous datasets, a challenge that remains central to AI pipelines today.
In large-scale AI systems, knowledge graphs—many of which adopt RDF or OWL-based schemas—serve as key instruments for data integration, enabling entity resolution, feature alignment, and graph-based feature engineering. Companies like Google, Amazon, and Microsoft have adopted internal ontologies to normalize concepts across business domains, facilitating model interoperability and multi-source training [66]. SPARQL’s graph-query paradigm also anticipates the emerging need for expressive retrieval interfaces in retrieval-augmented generation (RAG) systems.
Furthermore, recent approaches to embedding knowledge graphs into vector spaces for use in ML models draw on ontology axioms to constrain embedding learning, enabling logic-conformant representations [56]. Validation languages such as SHACL have also inspired schema-checking layers in AI pipelines, ensuring that training and inference data adhere to expected semantic contracts. In effect, these practices embed Semantic Web–style interoperability principles at the foundation of modern machine learning workflows.
Thus, although seldom cited by name, the Semantic Web’s influence persists through the design patterns, vocabularies, and validation logic that now underpin scalable and trustworthy AI data infrastructures.
@@ 3.6 Software Agents in the LLM Age 2017 onwards
The advent of large language models (LLMs) after the advent of the Transformer [74] has dramatically reshaped the landscape of software agent development, opening up new possibilities and challenges. Pre-LLM, agents often relied on meticulously crafted rules, ontologies, and knowledge bases, limiting their adaptability and generalizability [81, 64]. While these approaches provided a degree of control and explainability, they struggled to handle the complexity and ambiguity of real-world scenarios. LLMs, trained on massive datasets, offer a fundamentally different approach. They possess an unprecedented ability to understand and generate natural language, reason, and even exhibit some degree of common sense, albeit imperfectly [10, 60, 77]. This inherent capability allows LLM-powered agents to interact with users and environments in a much more natural and flexible way, reducing the need for extensive manual programming and domain-specific knowledge engineering.
The integration of LLMs with agent frameworks has led to the emergence of ”LLM-powered agents” or ”LLM agents,” capable of performing a wider range of tasks than their predecessors. The simple definition for an LLM Agent is an artificial entity with prompt specifications (initial state), conversation trace (state), and ability to interact with the environments such as tool usage (action) [11, 55]. These agents can leverage LLMs for various functions, including natural language understanding (NLU), natural language generation (NLG), planning, decision-making, and even tool use [84, 76, 85]. For instance, an LLM agent can be prompted to break down a complex task into smaller, manageable sub-tasks, search the web for relevant information, interact with APIs to execute actions, and summarize the results for the user [86, 67]. Frameworks like LangChain [12] and AutoGen [83] provide tools and abstractions to facilitate the development and deployment of these agents, enabling developers to chain together LLM calls, external tools, and custom logic. This marks a significant shift from symbolic AI to a more data-driven, emergent approach to agent development.
@@
The ongoing research in LLM-powered agents focuses on addressing these limitations and exploring new capabilities with the emerging multi-agent systems (MAS). This includes developing methods for improving the factual accuracy and reliability of LLMs, designing more robust and explainable agent architectures, and creating effective evaluation metrics to assess agent performance across diverse tasks [54, 35]. The convergence of LLMs and software agents represents a significant step towards building more general-purpose, adaptable, and intelligent systems, with the potential to transform how we interact with computers and automate complex tasks. However, responsible development and deployment, with a keen awareness of the inherent challenges, are crucial to realizing the full potential of this technology.
3.7 Prominent LLM Agent Frameworks and Libraries
A fast-maturing ecosystem of open-source frameworks now captures the canonical design patterns for agentic large-language-model (LLM) systems, letting researchers focus on reasoning strategies rather than boiler-plate glue code. LangChain supplies composable primitives—tools, memories and chains—that underpin many production chatbots, document-question-answering (QA) pipelines and research assistants. Building on the same abstractions, LangGraph introduces a graph-based execution engine with explicit state, branching and parallelism, which is invaluable for long-horizon planning and multi-agent coordination.
Microsoft’s AutoGen formalises “chat-loop” interactions in which several LLMs, optionally augmented by external tools or human feedback, iteratively critique and refine one another to solve non-trivial tasks. Where private or proprietary data are involved, LlamaIndex offers a complete data layer—index construction, retrieval interfaces and knowledge-graph support—for retrieval-augmented generation (RAG) over local corpora. CrewAI targets role-oriented workflows: agents endowed with complementary expertise share a common memory to tackle coordinated problem solving or simulation environments. For production use, Microsoft’s Semantic Kernel provides a language-agnostic plug-in system, a symbolic planner and robust security controls that simplify the integration of generative AI into enterprise software stacks. Google’s Agent Development Kit (ADK) implements the Agent-to-Agent (A2A) protocol and offers orchestration, safety guards and evaluation utilities for multi-agent research at scale. Also from Google, Genkit delivers a code-first, TypeScript workflow with flow orchestration, prompt/version control (.prompt files), observability and one-command Cloud Run deployment. Finally, Marvin provides a pythonic decorator-based API that turns ordinary functions into threaded, asynchronous “AI functions,” making rapid prototyping inside data-science or machine-learning pipelines straightforward. Additional libraries—including Haystack, Embedchain, SuperAGI, Dify and OpenAgents—continue to expand the design space and keep the agent-framework landscape highly dynamic.
3.8 Could Semantic Web Principles Complement LLMs?
Large-language-model pipelines still lack explicit semantics, provenance, and verifiability. Drawing on the Semantic-Web stack, we outline three concrete integration patterns that harness ontologies and knowledge graphs to mitigate hallucination and to provide user-facing rationales.
Why ontologies? Unlike the high-dimensional token spaces in which large language models (LLMs) operate, an ontology offers a canonical, machine-interpretable vocabulary of entities, relations and axioms. Mapping generated spans to URI-identified concepts therefore gives each symbol an unambiguous referent, providing the first step toward faithful explanations [14].
Complementary trade-offs. OWL and RDF guarantee sound, monotonic inference but require careful manual modelling, limit expressivity to stay decidable, adopt an Open World Assumption and incur worst-case exponential reasoning cost. LLMs sit at the opposite end of the spectrum: they learn broad coverage automatically, answer in closed-world style with constant-time decoding, and tolerate noisy data—yet offer no built-in notion of truth, identity or provenance. Combining the two thus pairs the symbolic rigour of OWL/RDF with the statistical reach of LLMs.
Pattern 1 – ontology-constrained decoding. During generation the model is prompted or fine-tuned so that every entity mention must be resolvable to an ontology concept; illegal continuations are pruned. This drastically reduces hallucinations and yields provenance for every token. A practical example is the plant-disease classifier of [2], where OWL reasoning is used online to check that the captioned visual attributes entail a unique disease class.
Pattern 2 – knowledge-graph RAG. Retrieval-augmented generation in which the retriever indexes a knowledge graph rather than free text grounds answers in triples whose semantics are explicit. Recent systems such as SubgraphRAG show that even small models achieve state-of-the-art accuracy and produce subgraph-level rationales that users can inspect [41].
Pattern 3 – neuro-symbolic post-hoc explanation. After an LLM produces an answer, a reasoning engine aligns the answer with the ontology, derives minimal entailment paths, and verbalises them as “because-clauses.” Confalonieri and Guizzardi identify reference modelling, commonsense reasoning and knowledge-refinement as the three main roles ontologies play in such explanatory pipelines [14].
3.9 How AI Has Shaped Real-World Semantic Web Applications
Since the publication of the 2001 Semantic Web vision, advances in AI have continually reshaped how Semantic Web technologies are applied in practice. While the original vision emphasized agent-based reasoning over RDF/OWL graphs, real-world adoption accelerated primarily in areas where AI provided scalable solutions for search, classification, and entity resolution.
Notably, the integration of AI with Semantic Web principles led to the widespread adoption of knowledge graphs, now core components of systems like Google’s Knowledge Panel and Facebook’s Entity Graph. These applications use statistical entity linking and machine learning–driven relation extraction, but rely on RDF-style schemas for consistent interpretation. Similarly, the emergence of schema.org as a lightweight ontology for Web content marked a convergence between SEO optimization and semantic markup, driven by AI-based search and ranking algorithms.
In recent years, machine learning pipelines have begun to adopt semantic validation layers—such as SHACL—for ensuring data integrity, especially in regulated domains like healthcare and finance. AI-driven recommendation systems now leverage ontologies to improve explainability and diversity of results, while retrieval-augmented generation (RAG) models increasingly use graph-based backends to ensure factual grounding.
Thus, while the original agent-centric vision remains largely aspirational, AI has incrementally transformed the Semantic Web into a practical substrate for large-scale, machine-interpretable data systems. These hybrid architectures demonstrate that AI advances have not only extended but redefined the scope of Semantic Web applications in real-world systems.